# Hydrodynamics of the electronic Fermi liquid: a pedagogical overview

### v1

@@ II.6 Thermal and thermoelectric transport
where $n_{E}$ is the energy density and $j_{E}$ is the energy current (energy flux density). The constitutive relations for the clean Fermi liquid can again be derived from the Boltzmann kinetic equation. For simplicity, we take the constitutive relations
@@
$\rho c_{p}\left(\partial_{t}T+u_{i}\partial_{i}T\right)=\kappa\partial^{2}T+\sigma^{\prime}_{ij}\partial_{i}u_{j}+\rho\gamma u^{2}$ (30)
@@
With the thermal transport equation [Eq. (30)] in hand, one can study the interplay between hydrodynamic flow and heat transport in materials. A number of works have studied thermoelectric effects in hydrodynamic systems using the Boltzmann formalism; the same techniques used to calculate viscosity can be used for thermal conductivity and the thermoelectric tensor. In general, hydrodynamic flow entails significant deviations from the Mott relation between thermopower and conductivity and the Wiedemann-Franz law between thermal and electrical conductivity (see Sec. II.7), which are hallmarks of the usual metallic Fermi liquid [114, 115, 116, 117, 27, 98, 118, 119, 120, 99, 121].
@@ II.7 Wiedemann-Franz Law
For a metallic Fermi liquid, both the electrical $\sigma$ and thermal $\kappa$ conductivities  [^8] The thermal conductivity here corresponds only to the electron contribution; phononic contributions are neglected. are controlled by a single time scale: the momentum relaxation time $\tau_{\text{mr}}$. As a result, the ratio
@@
is a universal constant at temperatures $T$ much smaller than the Fermi energy, where $e$ is the fundamental electric charge; $L_{\text{WF}}$ is known as the Lorenz number. Eq. (31) is known as the Wiedemann-Franz (WF) law. Significant departures from the WF law, therefore, signify non-(metallic)-Fermi liquid physics and can serve as supplementary evidence for hydrodynamic behavior.
@@
Before we discuss the WF law in the context of hydrodynamics, however, we need to deal with some subtleties. As previously discussed, the local conductivity $\sigma$ is an ill-defined quantity in hydrodynamics. There are therefore two ways to define the WF ratio. The first way is to ignore viscosity entirely and only take the ohmic contribution to conductivity, namely $\sigma=nq^{2}/(m\gamma_{\text{mr}})$. This treatment is equivalent to computing the WF ratio for uniform flow in an infinite domain, and it is a common definition in theory papers. The second way is to define the WF ratio with macroscopic conductances, namely $L_{\text{WF}}\equiv G_{\text{th}}/(G_{\text{el}}T)$, where $G_{\text{th}}$ and $G_{\text{el}}$ are the thermal and electrical conductances, respectively. This second definition includes viscous corrections to $G_{\text{el}}$ and corresponds to the typical experimental approach. While these two approaches agree in the ohmic regime, they are distinct in viscous hydrodynamics; $G_{\text{el}}$ depends on both the viscosity and the geometry of the sample. This dependence is is exemplified for the clean Fermi liquid with $\gamma_{\text{mr}}=0$; in a finite geometry with no-slip boundary conditions, the first definition diverges while the second remains finite.
@@
Regardless of the choice of definition, hydrodynamic systems violate the WF law. This violation has been computed theoretically [114, 127, 117, 118, 128] and observed experimentally [1, 47, 46, 51, 102]. For a Fermi liquid, the WF ratio is expected to be suppressed. The WF violation can be qualitatively understood as follows. Due to momentum conservation, electron-electron collisions do not degrade the flow of electron current, allowing the electrical conductance to be large. However, electron-electron collisions do not conserve the energy current (though they do conserve total energy) and thus they reduce the thermal conductivity. This decoupling of the electrical and thermal transport times can lead to a significant violation of the WF law in hydrodynamic materials, and in particular a suppression for a clean Fermi liquid. (As we discuss below in Sec. III.2, the opposite is true in the “Dirac fluid” regime, where electrons and holes are present in nearly equal number: electron-hole collisions degrade the electric current but not the energy current, so that $L_{\text{WF}}$ is substantially enhanced from the WF value.)
@@ II.8 Hydrodynamic flow in a smooth potential landscape
Thus far, we have only discussed two sources of momentum relaxation. The first is boundary scattering off of rigid walls. These are objects with hard boundaries that are large compared to the mean free path $\ell_{\text{ee}}$ for electron-electron collisions, and which give rise to frictional (e.g. no-slip) boundary conditions and effects like the Stokes paradox [129, 130, 58, 56, 57]. The second is momentum-relaxing scattering, e.g. from disorder. In this sense we have considered disorder only as a contributor to the finite momentum-relaxation rate $\gamma_{\text{mr}}$. Such a description is valid when the disorder potential that produces electron scattering varies over length scales that are much shorter than $\ell_{ee}$. In this section, we discuss a third situation: flow in a smoothly varying potential. In this situation, the current density adjusts itself continuously in response to the potential, and one cannot reduce the effects of the potential simply to a uniform momentum relaxation rate. Similarly, current flows “over” regions of varying potential, and the potential cannot be reduced to a set of “hard-wall” obstacles. Such smooth disorder potentials arise naturally in 2D electron systems due to stray charges embedded in the substrate or in a nearby delta-doping layer [131, 132]. Moiré potentials in twisted or lattice-mismatched 2D systems also provide a smooth potential, which is periodic rather than random [133]. We comment below on the distinction between periodic and random potentials in terms of their effect on hydrodynamic electrons.
In Ref. 122, the authors studied the resistivity associated with hydrodynamic electrons flowing through a smooth potential, and they pointed out that the energy dissipation associated with resistivity arises from two competing effects. First, viscous dissipation is proportional to the squared spatial gradient of the fluid velocity (i.e. $\propto\eta(\partial_{i}u_{j})^{2}$), so that viscous forces favor a spatially uniform flow profile. Second, disorder causes the equilibrium electron density to vary in space, which implies variations in the electronic entropy density. Current flowing across an entropy gradient produces an opposing voltage via the thermoelectric effect, which leads to energy dissipation. Thus, thermoelectric forces favor current to be strongly concentrated along equipotential contours, where the entropy density is constant. The optimal flow pattern can be thought of as that which minimizes the total power dissipated (or, more precisely, which minimizes the entropy generated within the sample [123]).
@@
where $\delta s$ is the typical spatial fluctuation in entropy density (at equilibrium) and $\xi$ is the spatial period of the potential. Ref. [123] calculated the numeric prefactors of this equation for different types of periodic potentials and pointed out that, if one assumes Fermi liquid scaling of the viscosity ($\eta\propto T^{-2}$), thermal conductivity ($\kappa\propto T^{-1}$), and entropy density ($\delta s\propto T$), then Eq. (32) implies a linear-in-$T$ behavior of the resistivity. This hydrodynamic flow mechanism is therefore a candidate explanation for the linear-in-$T$ resistivity observed in various strongly-correlated 2D moiré materials [134, 135, 136].
@@ II.9 Current noise of hydrodynamic electrons
Shot noise is associated with the discreteness of the charge carriers. To understand its origin in a slightly oversimplified way, consider a current $I$ flowing through a resistor. During a time interval $\Delta t$, this current is associated with a certain number $N=I\Delta t/q$ of charge carriers moving from the source to the drain electrodes. If the charge carriers move independently of each other at some finite rate, then there are statistical fluctuations in $N$ over any particular time interval, and these are of order $\Delta N\sim\sqrt{N}$; the arrivals are described by a Poisson process. Consequently, there are statistical fluctuations in the average current over that time interval of order $\Delta I=q\Delta N/\Delta t$. The noise power $(\Delta I)^{2}$ is therefore given by $q^{2}N/(\Delta t)$, or $(\Delta I)^{2}\sim|q|I\Delta f$, where $\Delta f\sim 1/\Delta t$ is the frequency bandwidth over which the current fluctuations are measured  [^9] The shot noise is sometimes quoted with an extra prefactor of $2$, i.e. $S(\omega)\equiv\langle I(\omega)I(0)\rangle=2eI$. This depends on the convention of Fourier transform; i.e. if one is computing the full two-sided Fourier transform (as done here) it does not have this factor of $2$. However, if one is considering the one-sided Fourier transform or spectral noise density (as is commonly done in the literature), one obtains an additional factor of $2$ Thus, shot noise as an experimental tool gives a direct measurement of the quasiparticle charge $q$ (see, e.g., [138, 139] for general reviews about shot noise). Shot noise measurements have been used, for example, to confirm the existence of fractionally charged quasiparticles in the fractional quantum Hall effect [140, 141].
@@
However, even when no dc current is flowing, there are still statistical fluctuations about $I=0$ that are associated with thermal fluctuations in the velocity of charge carriers. The corresponding current noise is called Johnson-Nyquist (JN) noise [142, 143]. The classical form of JN noise is $(\Delta I)^{2}=(2k_{B}T/R)\Delta f$, which can be thought of as one of the many incarnations of the fluctuation-dissipation theorem  [^10] Again, this is sometimes written with an additional factor of $2$ due to Fourier transform conventions. See footnote 9.. JN noise measurements are also a powerful experimental tool, since they provide a direct readout of the electron temperature (provided that one separately measures the two-terminal resistance $R$). Such noise thermometry can be exploited to make sensitive measurements of the electronic thermal conductivity and specific heat, which are accomplished by depositing a known amount of heat into the electron system and measuring its increase in temperature through JN noise [144, 145, 1, 146]. JN noise measurements also allow one to construct ultrasensitive light detectors (bolometers), in which the temperature increase of electrons that results after light absorption is detected through an increase in JN noise [147, 148, 149, 150, 151].
@@ II.10 Nonlinear effects
arising from the ratio of the momentum-relaxation term with the convective force, in analogy to the usual Reynolds number. For nonlinear effects to become important, both $\operatorname{Re}$ and $\operatorname{Re}_{\text{mr}}$ must be sufficiently large. In classical fluid dynamics, turbulent behavior begins around $\operatorname{Re}\sim 10^{3}$, though nonlinear effects can appear at much lower $\operatorname{Re}\sim 1$. Unfortunately, current experiments can only reach $\operatorname{Re},\operatorname{Re}_{\text{mr}}\sim 10^{-2}$, which makes probing nonlinear effects challenging.
@@
which is known as the Bernoulli equation. Despite its simplicity, the equation is nonlinear in $u$. For an incompressible flow where the inlet and outlet widths are not identical, e.g. a Venturi tube geometry, this nonlinearity gives rise to a non-linear and singular $I$-$V$ characteristic with $I\propto\sqrt{V}$ arising from convective acceleration [18]. This $I$-$V$ characteristic has a singular point, below which may exist a potential instability towards turbulent and/or intermittent flow. Typical nonlinear perturbative calculations of $I$ in powers of $V$ may have difficulty recovering this result due to the square-root singularity; the hydrodynamic approach, however, displays this nonlinearity straightforwardly.
@@
At sufficiently high Reynolds numbers, one can also induce turbulent behavior. Theoretical works have proposed to observe phenomena such as the Rayleigh-Bénard instability [126], the Kelvin-Helmholtz instability [170], and preturbulent behavior via current or electric potential fluctuations [171, 172, 173]. There has also been a proposal to observe the nonlinear dynamics of hot spot relaxation[55]. As previously mentioned, however, reaching the large drive parameters (e.g. large Reynolds numbers) necessary for these strongly nonlinear effects seems to be experimentally difficult [174, 5, 18].
@@ III.2 Electron-Hole and Dirac Fluid Hydrodynamics
At nonlinear order, however, the equations of motion are changed relative to the case of a parabolic dispersion. Due to the linear Dirac dispersion, the fluid is no longer Galilean-invariant, and this lack of invariance leads to corrections to the convective term in addition to other nonlinearities [202, 162, 55, 17, 163, 208, 5]. Predictions regarding modifications to plasma waves and oscillations have been made, though detailed experimental confirmation is still lacking. We remark that a recent experiment [209] has observed the relativistic sound mode (demon mode) with characteristic speed $v_{F}/\sqrt{2}$ in graphene [5, 55, 163, 210], though this effect arises from linear response.

### v2

@@ II.6 Thermal and thermoelectric transport
where $n_{E}$ is the energy density and $j_{E}$ is the energy current (energy flux density). The constitutive relations for the clean Fermi liquid can again be derived from the Boltzmann kinetic equation [71, 70, 72]. For simplicity, we take the constitutive relations
@@
$\rho c_{p}\left(\partial_{t}T+u_{i}\partial_{i}T\right)=\kappa\partial^{2}T+\sigma^{\prime}_{ij}\partial_{i}u_{j}+\rho\gamma_{\text{mr}}u^{2}$ (30)
@@
With the thermal transport equation [Eq. (30)] in hand, one can study the interplay between hydrodynamic flow and heat transport in materials. A number of works have studied thermoelectric effects in hydrodynamic systems using the Boltzmann formalism; the same techniques used to calculate viscosity can be used for thermal conductivity and the thermoelectric tensor. In general, hydrodynamic flow entails significant deviations from the Mott relation between thermopower and conductivity and the Wiedemann-Franz law between thermal and electrical conductivity [135, 136, 137, 138, 43, 119, 139, 140, 141, 120, 142] (see Sec. II.7), which are hallmarks of the usual metallic Fermi liquid [10].
@@ II.7 Wiedemann-Franz Law
For a metallic Fermi liquid, both the electrical $\sigma$ and thermal $\kappa$ conductivities  [^9] The thermal conductivity here corresponds only to the electron contribution; phononic contributions are neglected. are controlled by a single time scale: the momentum relaxation time $\tau_{\text{mr}}$ (see Eq. (16) with $I_{ee}\equiv 0$). As a result, the ratio
@@
is a universal constant at temperatures $T$ much smaller than the Fermi energy, where $e$ is the fundamental electric charge; $L_{\text{WF}}$ is known as the Lorenz number. Eq. (31) is known as the Wiedemann-Franz (WF) law, and is derived in the relaxation-time approximation [10]. Significant departures from the WF law, therefore, signify non-(metallic)-Fermi liquid physics and can serve as supplementary evidence for hydrodynamic behavior.
@@
Before we discuss the WF law in the context of hydrodynamics (i.e. turning on $I_{ee}\neq 0$), however, we need to deal with some subtleties. As previously discussed, the local conductivity $\sigma$ is an ill-defined quantity in hydrodynamics. There are therefore two ways to define the WF ratio. The first way is to ignore viscosity entirely and only take the ohmic contribution to conductivity, namely $\sigma=nq^{2}/(m\gamma_{\text{mr}})$. This treatment is equivalent to computing the WF ratio for uniform flow in an infinite domain, and it is a common definition in theory papers. The second way is to define the WF ratio with macroscopic conductances, namely $L_{\text{WF}}\equiv G_{\text{th}}/(G_{\text{el}}T)$, where $G_{\text{th}}$ and $G_{\text{el}}$ are the thermal and electrical conductances, respectively. This second definition includes viscous corrections to $G_{\text{el}}$ and corresponds to the typical experimental approach. While these two approaches agree in the ohmic regime, they are distinct in viscous hydrodynamics; $G_{\text{el}}$ depends on both the viscosity and the geometry of the sample. This dependence is is exemplified for the clean Fermi liquid with $\gamma_{\text{mr}}=0$; in a finite geometry with no-slip boundary conditions, the first definition diverges while the second remains finite.
@@
Regardless of the choice of definition, hydrodynamic systems violate the WF law. This violation has been computed theoretically [135, 148, 138, 139, 149] and observed experimentally [1, 63, 62, 67, 123] with up to order-of-magnitude violations. For a clean Fermi liquid, the WF ratio is expected to be suppressed. The WF violation can be qualitatively understood as follows. Due to momentum conservation, electron-electron collisions do not degrade the flow of electron current, allowing the electrical conductance to be large. However, electron-electron collisions do not conserve the energy current (though they do conserve total energy) and thus they reduce the thermal conductivity. This decoupling of the electrical and thermal transport times can lead to a significant violation of the WF law in hydrodynamic materials, and in particular a suppression for a clean Fermi liquid. This is sometimes handwavingly presented as $G_{\text{th}}\sim(\gamma_{\text{mr}}+\gamma_{\text{ee}})^{-1}$, but $G_{\text{el}}\sim\gamma_{\text{mr}}^{-1}$; we emphasize that this approximation, though helpful, should not be taken too literally due to various subtleties such as those described above. (As we discuss below in Sec. III.2, the opposite is true in the “Dirac fluid” regime, where electrons and holes are present in nearly equal number: electron-hole collisions degrade the electric current but not the energy current, so that $L_{\text{WF}}$ is substantially enhanced from the WF value.)
@@ II.8 Hydrodynamic flow in a smooth potential landscape
Thus far, we have only discussed two sources of momentum relaxation. The first is boundary scattering off of rigid walls. These are objects with hard boundaries that are large compared to the mean free path $\ell_{\text{ee}}$ for electron-electron collisions, and which give rise to frictional (e.g. no-slip) boundary conditions and effects like the Stokes paradox [150, 151, 79, 77, 78]. The second is momentum-relaxing scattering, e.g. from disorder. In this sense we have considered disorder only as a contributor to the finite momentum-relaxation rate $\gamma_{\text{mr}}$. Such a description is valid when the disorder potential that produces electron scattering varies over length scales that are much shorter than $\ell_{ee}$.
In this section, we discuss a third situation: flow in a smoothly varying potential. In this case, the current density adjusts itself continuously in response to the potential, and one cannot reduce the effects of the potential simply to a uniform momentum relaxation rate. Similarly, current flows “over” regions of varying potential, and the potential cannot be reduced to a set of “hard-wall” obstacles. Due to the slowly-varying nature of the potential, on the hydrodynamic coarse-graining scale $\ell_{ee}$ the potential is approximately constant; we remark that even a strong potential can be treated within the kinetic theory approach so long as the variation is slow with respect to $\ell_{ee}$. Such smooth disorder potentials arise naturally in 2D electron systems due to stray charges embedded in the substrate or in a nearby delta-doping layer [152, 153]. Moiré potentials in twisted or lattice-mismatched 2D systems also provide a smooth potential, which is periodic rather than random [154]. We comment below on the distinction between periodic and random potentials in terms of their effect on hydrodynamic electrons.
In Ref. 143, the authors studied the resistivity associated with hydrodynamic electrons flowing through a smooth potential, and they pointed out that the energy dissipation associated with resistivity arises from two competing effects. First, viscous dissipation is proportional to the squared spatial gradient of the fluid velocity (i.e. $\propto\eta(\partial_{i}u_{j})^{2}$), so that viscous forces favor a spatially uniform flow profile. Second, disorder causes the equilibrium electron density to vary in space, which implies variations in the electronic entropy density. Current flowing across an entropy gradient produces an opposing voltage via the thermoelectric effect, which leads to energy dissipation. Thus, thermoelectric forces favor current to be strongly concentrated along equipotential contours, where the entropy density is constant. The optimal flow pattern can be thought of as that which minimizes the total power dissipated (or, more precisely, which minimizes the entropy generated within the sample [144]).
@@
where $\delta s$ is the typical spatial fluctuation in entropy density (at equilibrium) and $\xi$ is the spatial period of the potential. Ref. [144] calculated the numeric prefactors of this equation for different types of periodic potentials and pointed out that, if one assumes Fermi liquid scaling of the viscosity ($\eta\propto T^{-2}$), thermal conductivity ($\kappa\propto T^{-1}$), and entropy density ($\delta s\propto T$)  [^10] We remark that the notation for $\delta s$ differs from that used in e.g. Ref. [155]. Here, $\delta s$ refers to the typical spatial variation of the entropy density as a result of the statically-imposed external potential, while in Ref. [155] it refers to fluctuations associated with oscillatory modes. As a result, the relation $\delta s\propto T$ is just the statement of how the entropy density of a Fermi liquid in static equilibrium changes with temperature., then Eq. (32) implies a linear-in-$T$ behavior of the resistivity. This hydrodynamic flow mechanism is therefore a candidate explanation for the linear-in-$T$ resistivity observed in various strongly-correlated 2D moiré materials [156, 157, 158].
@@ II.9 Current noise of hydrodynamic electrons
Shot noise is associated with the discreteness of the charge carriers. To understand its origin in a slightly oversimplified way, consider a current $I$ flowing through a resistor. During a time interval $\Delta t$, this current is associated with a certain number $N=I\Delta t/q$ of charge carriers moving from the source to the drain electrodes. If the charge carriers move independently of each other at some finite rate, then there are statistical fluctuations in $N$ over any particular time interval, and these are of order $\Delta N\sim\sqrt{N}$; the arrivals are described by a Poisson process. Consequently, there are statistical fluctuations in the average current over that time interval of order $\Delta I=q\Delta N/\Delta t$. The noise power $(\Delta I)^{2}$ is therefore given by $q^{2}N/(\Delta t)$, or $(\Delta I)^{2}\sim|q|I\Delta f$, where $\Delta f\sim 1/\Delta t$ is the frequency bandwidth over which the current fluctuations are measured  [^11] The shot noise is sometimes quoted with an extra prefactor of $2$, i.e. $S(\omega)\equiv\langle I(\omega)I(0)\rangle=2eI$. This depends on the convention of Fourier transform; i.e. if one is computing the full two-sided Fourier transform (as done here) it does not have this factor of $2$. However, if one is considering the one-sided Fourier transform or spectral noise density (as is commonly done in the literature), one obtains an additional factor of $2$. . Thus, shot noise as an experimental tool gives a direct measurement of the quasiparticle charge $q$ (see, e.g., [160, 161] for general reviews about shot noise). Shot noise measurements have been used, for example, to confirm the existence of fractionally charged quasiparticles in the fractional quantum Hall effect [162, 163].
@@
However, even when no dc current is flowing, there are still statistical fluctuations about $I=0$ that are associated with thermal fluctuations in the velocity of charge carriers. The corresponding current noise is called Johnson-Nyquist (JN) noise [164, 165]. The classical form of JN noise is $(\Delta I)^{2}=(2k_{B}T/R)\Delta f$, which can be thought of as one of the many incarnations of the fluctuation-dissipation theorem  [^12] Again, this is sometimes written with an additional factor of $2$ due to Fourier transform conventions. See footnote 11.. JN noise measurements are also a powerful experimental tool, since they provide a direct readout of the electron temperature (provided that one separately measures the two-terminal resistance $R$). Such noise thermometry can be exploited to make sensitive measurements of the electronic thermal conductivity and specific heat, which are accomplished by depositing a known amount of heat into the electron system and measuring its increase in temperature through JN noise [166, 167, 1, 168]. JN noise measurements also allow one to construct ultrasensitive light detectors (bolometers), in which the temperature increase of electrons that results after light absorption is detected through an increase in JN noise [169, 170, 171, 172, 173].
@@ II.10 Nonlinear effects
arising from the ratio of the momentum-relaxation term with the convective force, in analogy to the usual Reynolds number. For nonlinear effects to become important, both $\operatorname{Re}$ and $\operatorname{Re}_{\text{mr}}$ must be sufficiently large. In classical fluid dynamics, turbulent behavior begins around $\operatorname{Re}\sim 10^{3}$, though nonlinear effects can appear at much lower $\operatorname{Re}\sim 1$  [^13] We are being approximate and schematic here regarding Reynolds numbers; the exact transition point to turbulence depends on the specific geometry in question and can vary widely.. Unfortunately, current experiments can only reach $\operatorname{Re},\operatorname{Re}_{\text{mr}}\sim 10^{-2}$ before risking obliteration of the device from thermal heating. This makes probing nonlinear effects challenging.
@@
which is known as the Bernoulli equation. Despite its simplicity, the equation is nonlinear in $u$. For an incompressible flow where the inlet and outlet widths are not identical, e.g. a Venturi tube geometry, this nonlinearity gives rise to a non-linear and singular II-$V$ characteristic with $I\propto\sqrt{V}$ arising from convective acceleration [21]. This II-$V$ characteristic has a singular point, below which may exist a potential instability towards turbulent and/or intermittent flow. Typical nonlinear perturbative calculations of $I$ in powers of $V$ may have difficulty recovering this result due to the square-root singularity; the hydrodynamic approach, however, displays this nonlinearity straightforwardly.
@@
At sufficiently high Reynolds numbers, one can also induce turbulent behavior. By analogy to the classical fluid, one also expects Kolmogorov scaling with a direct energy cascade in 3D [19], and an inverse energy cascade in 2D [192]. However, the presence of finite $\tau_{\text{mr}}$ will cut off and potentially spoil the cascading behavior (see e.g. Ref. [5]). Theoretical works have proposed to observe phenomena such as the Rayleigh-Bénard instability [147], the Kelvin-Helmholtz instability [193], and preturbulent behavior via current or electric potential fluctuations [194, 195, 196]. There has also been a proposal to observe the nonlinear dynamics of hot spot relaxation[76]. As previously mentioned, however, reaching the large drive parameters (e.g. large Reynolds numbers) necessary for these strongly nonlinear effects seems to be experimentally difficult [197, 5, 21].
@@ III.2 Electron-Hole and Dirac Fluid Hydrodynamics
At nonlinear order, however, the equations of motion are changed relative to the case of a parabolic dispersion. Due to the linear Dirac dispersion, the fluid is no longer Galilean-invariant, and this lack of invariance leads to corrections to the convective term in addition to other nonlinearities [225, 184, 76, 20, 185, 231, 5]. Predictions regarding modifications to plasma waves and oscillations have been made, though detailed experimental confirmation is still lacking. We remark that a recent experiment [232] has observed the relativistic sound mode (demon mode or energy wave) with characteristic speed $v_{F}/\sqrt{2}$ in graphene [5, 76, 185, 233], though this effect arises from linear response. This mode is analogous to second sound in typical Fermi liquids (e.g. liquid He-3), which is also an energy wave of speed $v_{F}/\sqrt{d}$ for space dimension $d$.